# On-the-fly Surrogation for Complex Nonlinear Dynamics*

### v1

@@ IV-A1 Reconstruction of the true nonlinear system from the linear dynamics
For the simulations, we use the Matlab in-built variable-step solver ode45 with the default parameters and step size, and simulate for 14 seconds. Next, we replicate the simulations using the same settings, but this time we use (12) instead of the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$. For the integral in (12) we use the Matlab in-built function integral [^1] In the future we will explore the effects of different numerical integration schemes, as for the time being, we use this built-in function., which computes a numerical approximation of the integral based on the global adaptive quadrature approach from [28], with the respective absolute and relative tolerance settings set to $10^{-5}$ and $10^{-5}$. For these simulations, ode45 requested a maximum of 151 evaluations of (12) with a maximum computation time of 0.16 seconds [^2] On a 64 bit laptop PC with 2.10 GHz 8-core CPU and 64 GB RAM. The simulation results, shown in Fig. 1, illustrate that the solutions of $\Sigma_{\mathrm{vdp}}^{0.5}$ and its reconstruction with Corollary 2 are identical, supporting the results in Subsection III-A.
@@ IV-A2 On-the-fly surrogation of the Van der Pol oscillator
Second, we find a surrogate model of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the proposed on-the-fly surrogation method. For this, we define the following set of $N=9$ points:
@@
Next, we obtain the set $\mathcal{M}_{\mathrm{vdp}}^{1}$ by linearizing $\Sigma_{\mathrm{vdp}}^{0.5}$ at each point in $\Theta_{\mathrm{vdp}}^{1}$, and construct the set $\mathcal{D}_{\mathrm{vdp}}^{1}=\{\Theta_{\mathrm{vdp}}^{1},\mathcal{M}_{\mathrm{vdp}}^{1}\}$. Then, we obtain the surrogate model $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$ with Algorithms 1 and 2. For this, we choose the RBF $\phi$ given in (27) and the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\beta_{3}x_{2}+\beta_{4}u$ and the Matlab built-in numerical integration scheme integral. Furthermore, we obtain the width parameter $c$ by minimizing (30) with Bayesian optimization [29] as (30) is a non-smooth function, but other optimization methods suitable for non-smooth objectives can be used instead. Now, we compare the surrogate model $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$ with $\Sigma_{\mathrm{vdp}}^{0.5}$ with CT simulations, where the simulation and numerical integration settings are set as described in the previous experiment IV-A1. For these simulations, ode45 requested a maximum of 145 evaluations of (28b) with the highest computation time of 1.54 seconds. The results are shown in Fig. 2, where we observe that with only $N=9$ distinct linearization points, the obtained surrogate model $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$ can approximate the behaviour of $\Sigma_{\mathrm{vdp}}^{0.5}$.
Figure 2: State responses of the CT simulation of the Van der Pol oscillator with $\eta=0.5$, where ( ) is the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$ and (   ) is the surrogate model $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$ obtained with the data set $\mathcal{D}_{1}$.
@@ IV-A3 On-the-fly surrogation of the Van der Pol oscillator for a range of system parameter values
Third, we consider that the system parameter $\eta$ in  (31) lies within the range $\eta\in[0.3,0.6]$. To capture the system dynamics for the different values of $\eta$, we define the set $\mathcal{D}_{\mathrm{vdp}}^{2}$, which contains $N=68$ system observations contained within the admissible spaces defined as:
@@
Now, we construct the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ and compare in simulation with the true nonlinear system, using the same settings for the surrogate model and the simulations described in the second experiment IV-A2. In addition, for these simulations we set $\eta$ to different fixed values contained in (33) that are not included in the data set $\Theta_{\mathrm{vdp}}^{2}$. In this case, ode45 also requested a maximum of 145 evaluations of (28b), and the highest computation time was of 1.81 seconds. The simulation results are displayed in Fig. 3, where it is shown that the surrogate model $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ approximates the behaviour of the true nonlinear system for different values of $\eta$.
@@ IV-B Interconnection of MSD systems
Next, similar as in Subsection IV-A2, we obtain the surrogate model $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ with Algorithms 1 and 2, where $\mathcal{D}_{\mathrm{msd}}$, the RBF $\phi$ in (27), the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\dots\beta_{11}x_{10}+\beta_{12}u$, and the width parameter $c$ obtained by minimizing (30) with Bayesian optimization are the inputs to Algorithm 1. Then, we test the obtained surrogate by executing one thousand continuous-time simulations, where the states corresponding to the mass velocities, i.e., $\{x_{i}\}_{i=6}^{10}$ are initialized at zero, and the states corresponding to the mass positions, i.e., $\{x_{i}\}_{i=1}^{5}$ are initialized at random values drawn within the admissible space defined in (35) with a uniform distribution. Moreover, the external force $u$ is set to $u(t)=0.7\sin(2\pi t)$, and each simulation is executed for 8 seconds using ode45 with the default parameters and step size. For these simulations, ode45 requested an average of 390 evaluations of (28b) with an average computation time of 4 seconds. In Fig. 5
Figure 5: Simulated position (top) and velocity (bottom) trajectories of the $5$-th mass of the MSD interconnection, where ( ) and (   ) are obtained from the nonlinear system $\Sigma_{\mathrm{msd}}$ and the surrogate model $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, respectively.
@@
we display the trajectory of the last mass obtained from some simulation results, where it is observed that the surrogate model $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ approximates true nonlinear system $\Sigma_{\mathrm{msd}}$ with remarkable accuracy. Furthermore, we compute the root mean square error (RMSE) for each state of the simulated $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. the simulated $\Sigma_{\mathrm{msd}}$. Then, we compute the expected probability density distribution of the RMSE data based on a normal kernel function [30], and although there are some outliers with error values of up to $6\cdot 10^{-3}$, most of the error across all simulations is below $10^{-3}$, as shown shown in Fig. 6.

### v2

@@ IV-A1 Reconstruction of the true nonlinear system from the linear dynamics
For the simulations, we use the Matlab in-built variable-step solver ode45 with the default parameters and step size, and simulate for 14 seconds. Next, we replicate the simulations using the same settings, but this time we use (12) instead of the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$. For the integral in (12) we use the Matlab in-built function integral with the default tolerances, which computes a numerical approximation of the integral based on the global adaptive quadrature approach from [28]. For these simulations, ode45 requested an average of 21300 evaluations of the Jacobian of (31) with an average computation time of 0.16 seconds [^1] On a 64 bit laptop PC with 2.10 GHz 8-core CPU and 64 GB RAM. The simulation results, shown in Fig. 1, illustrate that the solutions of $\Sigma_{\mathrm{vdp}}^{0.5}$ and its reconstruction with Corollary 2 are identical, with an average root-mean-square error (RMSE) of $3.3\times 10^{-16}$, supporting the results in Subsection III-A.
@@ IV-A2 On-the-fly surrogation of the Van der Pol oscillator
Second, we find a surrogate of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the proposed on-the-fly surrogation method. For this, we define the following set of $N=9$ points:
@@
Next, we construct the set $\mathcal{M}_{\mathrm{vdp}}^{1}$ by linearizing $\Sigma_{\mathrm{vdp}}^{0.5}$ at each point in $\Theta_{\mathrm{vdp}}^{1}$ and define the set $\mathcal{D}_{\mathrm{vdp}}^{1}=\{\Theta_{\mathrm{vdp}}^{1},\mathcal{M}_{\mathrm{vdp}}^{1}\}$. Then, we construct the interpolant $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$ with Algorithm 1, with the RBF $\phi$ given in (27) and the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\beta_{3}x_{2}+\beta_{4}u$. The width parameter $c$ is determined by minimizing (30) with Bayesian optimization [29], since (30) is a non-smooth function. Note that other optimization methods suitable for non-smooth objectives can also be employed. Finally, we construct the surrogate representations $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$ and $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$ using $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$ and Algorithm 2 with two distinct numerical integration schemes. These two schemes are integral from Matlab, used for $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$, and the composite Simpson’s 3/8 rule (CS38) [30, Chapter 5.1] with 6 equidistant intervals, used for $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$.
Figure 2: State responses of the CT simulation of the Van der Pol oscillator with $\eta=0.5$, where ( ) is obtained with the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$, ( ) with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$ and (   ) with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$.
Now, we compare the surrogates w.r.t. $\Sigma_{\mathrm{vdp}}^{0.5}$ by CT simulations, where the simulation and numerical integration settings for integral are set as described in the previous experiment IV-A1. For the simulations with $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$, ode45 requested an average of 20400 evaluations of $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, leading to an average computation time and RMSE of 0.655 seconds and 0.1, respectively. For the simulations with $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$, ode45 requested an average of 816 evaluations of $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, leading to an average computation time and RMSE of 0.028 seconds and 0.1, respectively. The results are shown in Fig. 2, where we observe that with only $N=9$ distinct linearization points, the obtained surrogates can approximate the behaviour of $\Sigma_{\mathrm{vdp}}^{0.5}$. Moreover, the use of CS38 drastically reduces the computation time, while the effects in numerical accuracy are negligible.
@@ IV-A3 On-the-fly surrogation of the Van der Pol oscillator for a range of system parameter values
Third, we consider that the system parameter $\eta$ in (31) lies within the range $\eta\in[0.3,0.6]$. To capture the system dynamics for the different values of $\eta$, we define the set $\mathcal{D}_{\mathrm{vdp}}^{2}$, which contains $N=68$ system linearizations contained within the admissible spaces:
@@
Now, we construct the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ and compare in simulation with the true nonlinear system, using the same settings for the interpolant and the simulations described in the second experiment IV-A2, and CS38 with 6 equidistant intervals. In addition, for these simulations we set $\eta$ to different fixed values contained in (33) that are not included in $\Theta_{\mathrm{vdp}}^{2}$. In this case, ode45 also requested a maximum of 816 evaluations of the interpolant, leading to an average computation time and RMSE of 0.045 seconds and 0.784, respectively. These results are displayed in Fig. 3.
@@ IV-B Interconnection of MSD systems
Similar as in Subsection IV-A2, we construct the interpolant $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, where $\mathcal{D}_{\mathrm{msd}}$, the RBF $\phi$ in (27), the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\dots\beta_{11}x_{10}+\beta_{12}u$, and the width parameter $c$ obtained by minimizing (30) with Bayesian optimization are the inputs to Algorithm 1. Then, we formulate the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ according to Algorithm 2 using $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ and CS38 with 6 equidistant intervals.
Next, we test $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ by executing one thousand CT simulations, where the states corresponding to the mass velocities, i.e., $\{x_{i}\}_{i=6}^{10}$ are initialized at zero, and the states corresponding to the mass positions, i.e., $\{x_{i}\}_{i=1}^{5}$ are initialized at random values drawn within the admissible space defined in (35) with a uniform distribution. Moreover, the external force $u$ is set to $u(t)=0.7\sin(2\pi t)$, and each simulation is executed for 8 seconds using ode45 with the default parameters and step size. For these simulations, ode45 requested an average of 2352 evaluations of $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, leading to an average computation time of 0.15 seconds. In Fig. 5
Figure 5: Simulated position (top) and velocity (bottom) trajectories of the $5$-th mass of the MSD interconnection, where ( ) and (   ) are obtained from the nonlinear system $\Sigma_{\mathrm{msd}}$ and the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, respectively.
@@
we display the trajectory of the last mass obtained from some simulation results, where it is observed that the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ approximates true nonlinear system $\Sigma_{\mathrm{msd}}$ with remarkable accuracy. Furthermore, we compute the RMSE for each state of the simulated $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. the simulated $\Sigma_{\mathrm{msd}}$. Then, we compute the expected probability density distribution of the RMSE data based on a normal kernel function [31], and besides some outliers with error values of up to $6\cdot 10^{-3}$, most of the error across all simulations is below $10^{-3}$, as shown in Fig. 6.